Please help me solve the problem;
C=$2.50+1.25(h-1)
Please show your work so I can understand and do it on my own. Thank you so much!

Answer by stanbon(75887)

C=$2.50+1.25(h-1)
---------
Assuming you are asked to solve for "h"::
1.25(h-1) = C - 2.5
----
h-1 = (C-2.5)/1.25
-----
h = (C-2.5)/1.25 + 1
------------------------
Cheers,
Stan H.
-----------------